A manufacturer of chocolate chips would like to know whether its bag filling machine works correctly at the 439.0 gram setting. It is believed that the machine is underfilling the bags. A 41 bag sample had a mean of 433.0 grams. A level of significance of 0.1 will be used. Determine the decision rule. Assume the variance is known to be 324.00. Enter the decision rule: Reject Ho if z<____ ????
Answer #1

The null and alternative hypotheses are

H0: 11 433 grams

H1: μ < 433 grams

For level of significance of 0.1, the critical value of z for left tail test is, -2.326

So, the decision rule is

Reject Ho if z< -2.326

Given, sigma^2 = 324

=> sigma = sqrt{324} = 18

Standard error of sample mean = o/v/n 18/V41 2.811128

Test statistic, z = (433 - 439) / 2.811128 = -2.134

As, the test statistic (-2.134) is not less than the critical z value (-2.326), we fail to reject H0 and conclude that there is no significant evidence that μ < 433 and the machine is underfilling the bags